How many tetracycline tablets with a an average mass of 115 mg can be made from 1.7 kg of tetracycline ?
1
Expert's answer
2018-12-17T06:38:10-0500

Answer:

1)Convert 1.7 kg into mg: 1.7 kg = 1.7 x 10^3 g = 1.7x 10^6 mg

2) Number of tetracycline tablets = 1.7 x 10^6 mg / 115 mg = 0.01478 x 10^6 = 14780